# FinGPT: Enhancing Sentiment-Based Stock Movement Prediction with Dissemination-Aware and Context-Enriched LLMs

### v1

@@ Introduction
This section outlines our data-driven methodologies. Specifically, we 1) increase stock price granularity and implemented news clustering in Data Processing part, and 2) incorporate contextual and more targeted instructions in Prompt Engineering part.
@@
The baseline method uses only weekly aggregate stock price movements (e.g., 3% weekly gain). To enhance prediction performance, we increase data granularity by incorporating daily closing prices and corresponding returns throughout each week. This granular approach serves two key purposes: 1) it reduces calculation uncertainty within the LLMs by providing explicit daily price movements, and 2) enables precise temporal alignment between price changes and news events, providing a basis for the differentiation of short-term and long-term impacts. We will refer to this method as HG in the following part.
@@
We instruct LLMs to differentiate between short-term and long-term impacts of news, as daily stock prices and returns can reveal immediate market reactions. This distinction is crucial because the influence of short-term news is often already reflected in the stock movements within the same week. For prompt template, see Figure 4.
@@
Built upon HG, we construct the news component using selected representative articles and their associated metadata (topic size and temporal coverage). Then, we enhance the instructions by describing the news component and providing guidelines for analyzing the impact of news dissemination on stock movement. For prompt template, see Figure 5.
@@
The training dataset pairs our structured input prompts (including company introduction, historical stock prices, related news, company fundamentals, and instructions for utilizing sentiment analysis for prediction) with GPT-4o generated analysis based on known future movements. Then, removing the ground truth stock price in the prompt, we use this dataset to fine-tune Llama3-8B for weekly stock movement prediction and evaluate both numerical accuracy and reasoning quality (see Appendix A for details.)

### v2

@@ Introduction
This section outlines our data-driven methodologies. Specifically, we 1) increase stock price granularity and implemented news clustering in the Data Processing part, and 2) incorporate contextual and more targeted instructions in Prompt Engineering part.
@@
The baseline method uses only weekly aggregate stock price movements (e.g. 3% weekly gain). To enhance prediction performance, we increase data granularity by incorporating daily closing prices and corresponding returns throughout each week. This granular approach serves two key purposes: 1) it reduces the uncertainty of the calculation within the LLMs by providing explicit daily price movements and 2) enables precise temporal alignment between price changes and news events, providing a basis for the differentiation of short- long-term impacts. We will refer to this method as HG in the following part.
@@
We instruct LLMs to differentiate between short-term and long-term impacts of news, as daily stock prices and returns can reveal immediate market reactions. This distinction is crucial because the influence of short-term news is often already reflected in stock movements within the same week. For a prompt template, see Figure 4.
@@
Built upon HG, we construct the news component using selected representative articles and their associated metadata (topic size and temporal coverage). Then, we enhance the instructions by describing the news component and providing guidelines for analyzing the impact of news dissemination on stock movement. For a prompt template, see Figure 5.
@@
The training dataset pairs our structured input prompts (including company introduction, historical stock prices, related news, company fundamentals, and instructions for utilizing sentiment analysis for prediction) with GPT-4o-generated analysis based on known future movements. Then, removing the ground truth stock price in the prompt, we use this dataset to fine-tune Llama3-8B for weekly stock movement prediction and evaluate both numerical accuracy and reasoning quality (see Appendix A for details.)